# Key Words.: stars: abundances – Local Group – galaxies: dwarf – galaxies: formation1 Introduction

### v1

@@
We present our analysis of VLT/UVES and X-shooter observations of six very metal-poor stars, including four stars at in the Fornax and Carina dwarf spheroidal galaxies (dSph). Until now, this metallicity range in these two galaxies was either hardly or not yet investigated. The chemical abundances of 25 elements are presented, based on 1D/LTE model atmospheres. We discuss the different elemental groups, and find that - and iron-peak elements in these two systems are generally in good agreement with the Milky Way halo at the same metallicity. Our analysis reveals that none of the six stars we studied exhibit carbon enhancement, which is noteworthy given the prevalence of carbon-enhanced metal-poor (CEMP-no; no Ba enhancement) stars in the Galaxy at similarly low metallicities. Our compilation of literature data shows that the fraction of CEMP-no stars in dSphs is significantly lower than in the Milky Way, and than in ultra faint dwarf galaxies. Furthermore, we report the discovery of the lowest metallicity, [Fe/H] = , r-process rich (r-I) star in a dSph galaxy. This star, fnx_06_019, has , and also shows enhancement of La, Nd, and Dy, . Our new data in Carina and Fornax help to populate the extremely low metallicity range in dSph galaxies, and add onto the evidence of a low fraction of CEMP-no stars in these systems.
Extremely metal-poor stars in the Fornax and Carina dwarf spheroidal galaxies††thanks: Based on UVES and X–Shooter observations collected at the ESO, programme ID 0100.D–0820 and 094.D–0853.
@@
We aim at understanding the characteristics of the first stars formed in the universe, from their imprints on low-mass stars in dwarf spheroidal galaxies (dSphs). The stellar abundance trends and dispersion of the most metal-poor stars reveal the nature of the now disappeared first stellar generation (e.g., mass, numbers), and the level of homogeneity of the primitive interstellar medium (ISM), e.g., size/mass of star-forming regions, as well as the nature and energy of the first supernovae (e.g. Koutsouridou et al., 2023). The proximity of the Local Group dSphs allows to derive chemical abundances in individual stars at comparable quality as in the Milky Way (MW). The confrontation of galaxies with very different evolutionary paths brings crucial information on the universality of the star-formation processes and chemical enrichment.
@@
Carina, Sextans, Sculptor, and Fornax are four Local Group dSphs which have so far triggered strong observational efforts from the galactic archaeology community. They provided the first evidence for distinct star formation histories and chemical evolution from those of the Milky Way at (e.g. Tolstoy et al., 2009). These four dwarf galaxies form a sequence of mass from the lowest to the highest mass limits of the classical dSphs. They followed very different evolution: while the majority of stars in Sextans and Sculptor formed during the first 4 Gyr (de Boer et al., 2012, 2014; Bettinelli et al., 2018, 2019), Carina is famous for its distinct star-forming episodes, which are separated by long periods of quiescence (de Boer et al., 2014). With a significantly larger stellar mass, Fornax has an extended star formation history, and a population dominated by intermediate age stars (de Boer et al., 2012). Furthermore, in Fornax there are also six known globular clusters (GCs; Hodge, 1961; Pace et al., 2021). The diversity of these four dSphs allows us to probe the relation between the very early stages of star formation and the subsequent evolutionary paths.
@@
While we now broadly understand their later stages of evolution, their earliest times are still poorly understood. So far, Sculptor and Sextans are the classical dSphs with the largest number of metal-poor stars observed at sufficient spectral quality to derive accurate chemical abundances. To date the most studied dwarf galaxy at low [Fe/H] is Sculptor, with 13 chemically analyzed metal-poor stars at [Fe/H] . (Tafelmeyer et al., 2010; Frebel et al., 2010; Starkenburg et al., 2013; Jablonka et al., 2015; Simon et al., 2015; Skúladóttir et al., 2021; Skúladóttir et al., 2023b). Only four EMP stars out of 14 known with have been studied in Sextans (Shetrone et al., 2001; Aoki et al., 2009; Starkenburg et al., 2013; Lucchesi et al., 2020; Theler et al., 2020). Fornax has only one known EMP star that has been chemically characterized, with total of 2 stars at [Fe/H] (Tafelmeyer et al., 2010; Lemasle et al., 2014), and Carina has 9 stars with [Fe/H] have which been chemically analysed (Koch et al., 2008; Venn et al., 2012; Susmitha et al., 2017; Norris et al., 2017; Hansen et al., 2023).
ID | Setting | range | S/N | V | ||
h:mn:s | :’:” | Å | [/pix] | [km s] | ||
@@
In the recent years it has become apparent that different dwarf galaxies show different abundance trends at low . Around , most dSph galaxies, e.g. Fornax, Sextans, and Sculptor (Tafelmeyer et al., 2010; Theler et al., 2020; Lucchesi et al., 2020; Skúladóttir et al., 2023b), present a plateau of enhanced [/Fe]+0.4 with a low scatter, similar to the Milky Way and the small ultra-faint dwarf galaxies (UFDs). However, there are some deviations from this, e.g. in Sculptor, where individual unusual low- stars have been detected at (e.g. Jablonka et al., 2015), and the [/Fe] plateau seems to break down towards the lowest (e.g. Skúladóttir et al., 2023b). In addition, dSphs and UFDs show significant differences in their Sr and Ba abundances. Typically the [Sr/Fe] and [Ba/Fe] in dSphs increase towards higher [Fe/H], e.g. in Sextans where at , and at (Lucchesi et al., 2020). On the contrary, [Sr/Fe] and [Ba/Fe] typically stay sub-solar at all [Fe/H] in UFDs (Ji et al., 2019). The dSph galaxies show an increasing trend in [Sr/Ba] towards the lowest . This is not seen in UFDs, and the Sextans and Ursa Minor dSphs seem to be at the lower mass limit of where this trend is visible (Mashonkina et al., 2017b; Reichert et al., 2020; Lucchesi et al., 2020). It is likely that this relates to the earliest production sites of Sr and Ba, however, these are still being debated.
The nucleosynthetic sites of the neutron-capture elements are still largely debated in the literature, in particular in regards to the rapid -process which occurs under high neutron flux (e.g. Schatz et al., 2022). Investigating stars in galaxies covering a wide mass range, provides very crucial constraints, and supports a double origin of -process elements (Skúladóttir & Salvadori, 2020), by a rare type of massive stars (e.g., Winteler et al., 2012), as well as neutron star mergers (e.g., Wanajo et al., 2014). Both in the Milky Way and dwarf galaxies, individual stars have been found to exhibit enhancement of r-process products. These stars are defined as (Beers & Christlieb, 2005): i) r-I stars when , ; ii) r-II stars when [Eu/Fe]+1; [Ba/Eu]0. From the survey carried out by Barklem et al. (2005), the expected frequencies of r-II and r-I stars in the Milky Way halo are 3% and 15%, respectively. In Fornax, 3 r-II stars have been identified at high (Reichert et al., 2021), and several r-I stars have been reported at in Letarte et al. (2010). In the Ursa Minor dSph, r-II stars have also been identified at (Shetrone et al., 2001; Aoki et al., 2007b). In addition, some UFDs have been found to host very r-rich stars (Ji et al., 2016; Hansen et al., 2017). These metal-poor r-rich stars are extremely important, since they likely show the imprints of a singular r-event. Understanding the distribution of such stars in different systems, therefore allows us to trace back the origin of the r-process enrichment.
@@ 2.1 Target selection, observations, and data reduction
The EMP candidates of this work are red giant branch (RGB) stars in the Fornax and Carina dSph galaxies (Fig. 1 and 2). Their selection was based on an estimation of their metallicity via the Calcium II triplet (CaT), [Fe/H] –2.5. Starkenburg et al. (2010) delivered a CaT calibration down to , which was applied to the samples of Battaglia et al. (2006) for Fornax and Koch et al. (2006) for Carina. The observational journal of the programme stars is in Table 1.
The two EMP candidates in Fornax, fnx_06_019 and fnx0579x1, were sufficiently bright to enable follow-up at high resolution with the UVES spectrograph (Dekker et al., 2000) mounted at the ESO VLT (programme ID 0100.D-0820(A)). We used dichroic 1 with the CCD#2 centered at 3900 Å and the CCD#3 centered at 5800 Å. Setting the slit width at 1.2″ led to a nominal resolution of R34,000. The full wavelength coverage is 3200–6800 Å (see Table 1 for details), and the effective usable spectral information starts from 3800 Å due to the lower SNR toward the blue part of the spectrum. Each star has been observed for a total of five hours, split into six individual sub-exposures.
The four EMP candidates in Carina, LG04c_0008, Car1_t200, Car1_t174, and Car1_t194 were observed (programme ID 094.D–0853(B)) with X-Shooter (Vernet et al., 2011). The UVB slit was open to 0.811 arcsec while the VIS slit was open to 0.911 arcsec, which led to a nominal resolution of R6,200 and R7,400 respectively. The total exposure time, in STARE mode, was 2.5 hours for LG04c_0008 and 3 hours for the other stars divided in 3 and 4 OBs of 3000 s, respectively. The usable wavelength range spans 3040–6800 Å.
In all cases, the reduced data, including bias subtraction, flat fielding, wavelength calibration, spectral extraction, and order merging, were taken from the ESO Science Archive Facility.
Table 1 provides the coordinates of our targets, the signal-to-noise ratios (SN), and the radial velocities as measured in each wavelength interval. Table 2 lists the optical and near-infrared magnitudes of our sample. Fig. 1 shows the Colour Magnitude Diagram (CMD) of the targets, and Fig. 2 indicates the spatial location of our targets relative to the position of other spectroscopic studies in Carina and Fornax.
@@ 2.2 Radial velocity measurements and normalization
The stellar heliocentric radial velocities (RVs) were measured with the IRAF111Image Reduction and Analysis Facility; Astronomical Source Code Library ascl:9911.002 task rvidlines on each individual exposure. The final RV is the average of these individual values weighted by their uncertainties. This approach allows us to detect possible binary stars, at least those whose RV variations can be detected within about one year. We did not find any evidence for binarity. After they were corrected for RV shifts, the individual exposures were combined into a single spectrum using the IRAF task scombine with a 2 to 3 sigma clipping. As a final step, each spectrum was visually examined, and the few remaining cosmic rays were removed with the splot routine.
The mean RV of each Fornax star (Table 1) coincides with the RV of Fornax, 54.1 0.5 , within the velocity dispersion , measured by Battaglia et al. (2006), and 55.46 0.63 , within the velocity dispersion , measured by Hendricks et al. (2014b). Similarly, the RVs of the Carina stars fall within the mean of the galaxy, 224.4 5.95 , measured by Lemasle et al. (2012); and 223.9 with =7.5 measured by Koch et al. (2006). This confirms that our stars are galaxy members. Spectra were normalized using DAOSPEC (Stetson & Pancino, 2008) for each of the wavelength ranges presented in Table 1. We used a 20 to 40 degree polynomial fit.
@@ 3.1 Line list and model atmospheres
We adopted the new MARCS 1D atmosphere models and selected the Standard composition class, that is, we included the classical -enhancement of +0.4 dex at low metallicity. They were downloaded from the MARCS web site (Gustafsson et al., 2008), and interpolated using Thomas Masseron’s code, which is available on the same web site222http://marcs.astro.uu.se. Inside a cube of eight reference models, this code performs a linear interpolation on three given parameters: T, g, and [Fe/H].
@@ 3.2 Photometric temperature and gravity
The atmospheric parameters were initially determined using photometric information as reported in Table 2. The first approximated determination of the stellar effective temperature, T, was based on the VI, VJ, VH, and VK color indices measured by Battaglia et al. (2006), and J and Ks photometry was taken from the VISTA commissioning data, which were also calibrated onto the 2MASS photometric system. We assumed (Cardelli et al., 1989) and = 0.03 for Fornax (Letarte et al., 2010) and = 0.061 for Carina (de Boer et al., 2014) for the reddening correction, respectively. The adopted photometric effective temperatures, , are listed in Table 3. They correspond to the simple average of the four color temperatures derived from , , , and with the calibration of Ramírez & Meléndez (2005).
@@
Because only very few Fe ii lines can be detected in our X-Shooter spectra, the determination of surface gravities from the ionisation balance of Fe I vs. Fe II was not possible. Non-local thermodynamic equilibrium (NLTE) effects also play a role at extremely low metallicity and impact the abundances of Fe I, with (Fe II–Fe I) up to +0.20 dex at [Fe/H] = (Mashonkina et al., 2017a), thus surface gravities were determined from their relation with :
@@
assuming = 4.44, = 5790 K, and M = 4.75 for the Sun. We adopted a stellar mass of 0.8 M and calculated the bolometric corrections using the Alonso et al. (1999) calibration, with a distance of d=138 kpc (Battaglia et al., 2006) for Fornax and d=106 kpc (de Boer et al., 2014) for Carina.
3.3 Final stellar parameters and abundance determination

### v2

@@
Extremely metal-poor stars in the Fornax and Carina dwarf spheroidal galaxies††thanks: Based on UVES and X-Shooter observations collected at the ESO, program ID 0100.D–0820 and 094.D–0853.
We present our analysis of VLT/UVES and X-shooter observations of six very metal-poor stars, including four stars at [Fe/H] in the Fornax and Carina dwarf spheroidal (dSph) galaxies. To date, this metallicity range in these two galaxies has not yet been investigated fully, or at all in some cases. The chemical abundances of 25 elements are presented, based on 1D and local thermodynamic equilibrium (LTE) model atmospheres. We discuss the different elemental groups, and find that - and iron-peak elements in these two systems are generally in good agreement with the Milky Way halo at the same metallicity. Our analysis reveals that none of the six stars we studied exhibits carbon enhancement, which is noteworthy given the prevalence of carbon-enhanced metal-poor stars without s-process enhancement (CEMP-no) in the Galaxy at similarly low metallicities. Our compilation of literature data shows that the fraction of CEMP-no stars in dSph galaxies is significantly lower than in the Milky Way, and than in ultra-faint dwarf galaxies. Furthermore, we report the discovery of the lowest metallicity, [Fe/H] = , r-process rich (r-I) star in a dSph galaxy. This star, fnx_06_019, has , and also shows enhancement of La, Nd, and Dy, . Our new data in Carina and Fornax help populate the extremely low metallicity range in dSph galaxies, and add to the evidence of a low fraction of CEMP-no stars in these systems.
@@
Our aim is to understand the characteristics of the first stars formed in the universe from their imprints on low-mass stars in dwarf spheroidal (dSph) galaxies. The stellar abundance trends and dispersion of the most metal-poor stars reveal the nature of the now disappeared first stellar generation (e.g., mass, numbers), and the level of homogeneity of the primitive interstellar medium (ISM) (e.g., size and/or mass of star-forming regions), as well as the nature and energy of the first supernovae (e.g., Koutsouridou et al., 2023). The proximity of the Local Group dSph galaxies allows the chemical abundances to be derived in individual stars at comparable quality to the Milky Way (MW). The confrontation of galaxies with very different evolutionary paths brings crucial information on the universality of the star formation processes and chemical enrichment.
@@
Carina, Sextans, Sculptor, and Fornax are four Local Group dSph galaxies that have triggered strong observational efforts from the galactic archaeology community. They provide the first evidence for star formation histories and chemical evolution distinct from those of the Milky Way at (e.g., Tolstoy et al., 2009). These four dwarf galaxies form a sequence of mass from the lowest to the highest mass limits of the classical dSph galaxies. They followed very different evolutions: while the majority of stars in Sextans and Sculptor formed during the first 4 Gyr (de Boer et al., 2012, 2014; Bettinelli et al., 2018, 2019), Carina is famous for its distinct star-forming episodes, which are separated by long periods of quiescence (de Boer et al., 2014). With a significantly higher stellar mass, Fornax has an extended star formation history and a population dominated by intermediate-age stars (de Boer et al., 2012). Furthermore, in Fornax there are also six known globular clusters (GCs; Hodge, 1961; Pace et al., 2021). The diversity of these four dSph galaxies allows us to probe the relation between the very early stages of star formation and the subsequent evolutionary paths.
@@
While we now broadly understand their later stages of evolution, their earliest times are still poorly understood. Sculptor and Sextans are the classical dSph galaxies with the largest number of metal-poor stars observed at sufficient spectral quality to derive accurate chemical abundances. To date the most studied dwarf galaxy at low [Fe/H] is Sculptor, with 13 chemically analyzed metal-poor stars at [Fe/H] . (Tafelmeyer et al., 2010; Frebel et al., 2010; Starkenburg et al., 2013; Jablonka et al., 2015; Simon et al., 2015; Skúladóttir et al., 2021, 2023b). Only 4 extremely metal-poor (EMP) stars out of 14 known with have been studied in Sextans (Shetrone et al., 2001; Aoki et al., 2009; Starkenburg et al., 2013; Lucchesi et al., 2020; Theler et al., 2020). Fornax has only one known EMP star that has been chemically characterized, with total of two stars at [Fe/H] (Tafelmeyer et al., 2010; Lemasle et al., 2014). Carina has nine stars with [Fe/H] that have been chemically analyzed (Koch et al., 2008; Venn et al., 2012; Susmitha et al., 2017; Norris et al., 2017; Hansen et al., 2023).
| ID | Setting | range | S/N | Vrad,helio | ||
| h:mn:s | :’:” | Å | [/pix] | [km s-1] | ||
@@
In recent years it has become apparent that different dwarf galaxies show different abundance trends at low . Around , most dSph galaxies (e.g., Fornax, Sextans, and Sculptor) (Tafelmeyer et al., 2010; Theler et al., 2020; Lucchesi et al., 2020; Skúladóttir et al., 2023b) present a plateau of enhanced [/Fe]+0.4 with a low scatter, similar to the Milky Way and the small ultra-faint dwarf galaxies (UFDs). However, there are some deviations from this, for example in Sculptor, where individual unusual low- stars have been detected at (e.g., Jablonka et al., 2015), and the [/Fe] plateau seems to break down toward the lowest (e.g., Skúladóttir et al., 2023b). In addition, dSph and UFD galaxies show significant differences in their Sr and Ba abundances. Typically, the values of [Sr/Fe] and [Ba/Fe] in dSph galaxies increase toward higher [Fe/H], for example in Sextans where at , and at (Lucchesi et al., 2020). On the contrary, [Sr/Fe] and [Ba/Fe] typically remain subsolar at all [Fe/H] in UFDs (Ji et al., 2019). The dSph galaxies show an increasing trend in [Sr/Ba] toward the lowest . This is not seen in UFDs, and the Sextans and Ursa Minor dSph galaxies seem to be at the lower mass limit of where this trend is visible (Mashonkina et al., 2017b; Reichert et al., 2020; Lucchesi et al., 2020). It is likely that this relates to the earliest production sites of Sr and Ba; however, this is still under debated.
The nucleosynthetic sites of the neutron-capture elements are still largely debated in the literature, in particular with regard to the rapid -process, which occurs under high neutron flux (e.g., Schatz et al., 2022). Investigating stars in galaxies covering a wide mass range provides very crucial constraints, and supports a double origin of -process elements (Skúladóttir & Salvadori, 2020) by a rare type of massive stars (e.g., Winteler et al., 2012), as well as neutron star mergers (e.g., Wanajo et al., 2014). In the Milky Way and and in dwarf galaxies individual stars have been found to exhibit enhancement of r-process products. These stars are defined as r-I stars when , and as r-II stars when [Eu/Fe]+1, [Ba/Eu]0 (Beers & Christlieb, 2005). From the survey carried out by Barklem et al. (2005), the expected frequency of r-II and r-I stars in the Milky Way halo is 3% and 15%, respectively. In Fornax, three r-II stars have been identified at high (Reichert et al., 2021), and several r-I stars have been reported at (Letarte et al., 2010). In the Ursa Minor dSph galaxy, r-II stars have also been identified at (Shetrone et al., 2001; Aoki et al., 2007b). In addition, some UFDs have been found to host very r-rich stars (Ji et al., 2016; Hansen et al., 2017). These metal-poor r-rich stars are extremely important since they likely show the imprints of a singular r-event. Understanding the distribution of such stars in different systems therefore allows us to trace back the origin of the r-process enrichment.
@@ 2.1 Target selection, observations, and data reduction
The EMP candidates in this work are red giant branch (RGB) stars in the Fornax and Carina dSph galaxies (Figs. 1 and 2). They were selected based on an estimation of their metallicity via the calcium II triplet (CaT), [Fe/H] –2.5. Starkenburg et al. (2010) delivered a CaT calibration down to , which was applied to the Battaglia et al. (2006) sample for Fornax and the Koch et al. (2006) sample for Carina. The observational journal of the program stars is in Table 1.
The two EMP candidates in Fornax, fnx_06_019 and fnx0579x1, were sufficiently bright to enable follow-up at high resolution with the UVES spectrograph (Dekker et al., 2000) mounted at the ESO VLT (program ID 0100.D-0820(A)). We used dichroic 1 with the CCD#2 centered at 3900 Å and the CCD#3 centered at 5800 Å. Setting the slit width at 1.2″ led to a nominal resolution of R34,000. The full wavelength coverage is 3200–6800 Å (see Table 1 for details), and the effective usable spectral information starts from 3800 Å due to the lower signal-to-noise ratio (S/N) toward the blue part of the spectrum. Each star was observed for a total of five hours, split into six individual sub-exposures. The final S/N ranges from 12 (blue part) to 45 (red part).
The four EMP candidates in Carina (LG04c_0008, Car1_t200, Car1_t174, and Car1_t194) were observed (program ID 094.D–0853(B)) with X-Shooter (Vernet et al., 2011). The UVB slit was open to 0.811 arcsec2, while the VIS slit was open to 0.911 arcsec2, which led to a nominal resolution of R6,200 and R7,400, respectively. The total exposure time, in STARE mode, was 2.5 hours for LG04c_0008 and 3 hours for the other stars divided in 3 and 4 OBs of 3000 s, respectively. The usable wavelength range spans 3040–6800 Å with S/N ranging from 24 to 40.
In all cases the reduced data, including bias subtraction, flat fielding, wavelength calibration, spectral extraction, and order merging, were taken from the ESO Science Archive Facility. Table 1 provides the coordinates of our targets, the signal-to-noise ratios, and the radial velocities as measured in each wavelength interval. Table 2 lists the optical and near-infrared magnitudes of our sample. Figure 1 shows the color–magnitude diagram (CMD) of the targets, and Fig. 2 indicates the spatial location of our targets relative to the position of other spectroscopic studies in Carina and Fornax.
@@ 2.2 Radial velocity measurements and normalization
The stellar heliocentric radial velocities (RVs) were measured with the IRAF111Image Reduction and Analysis Facility; Astronomical Source Code Library ascl:9911.002 task rvidlines on each individual exposure. The final RV is the average of these individual values weighted by their uncertainties. This approach allows us to detect possible binary stars, at least those whose RV variations can be detected within about one year. We did not find any evidence for binarity. After they were corrected for RV shifts, the individual exposures were combined into a single spectrum using the IRAF task scombine with a 2–3 clipping. As a final step, each spectrum was visually examined, and the few remaining cosmic rays were removed with the splot routine.
The mean RV of each Fornax star (Table 1) coincides with the RV of Fornax, 54.1 0.5 , within the velocity dispersion measured by Battaglia et al. (2006), and 55.46 0.63 , within the velocity dispersion measured by Hendricks et al. (2014b). Similarly, the RVs of the Carina stars fall within the mean of the galaxy, 224.4 5.95 measured by Lemasle et al. (2012), and 223.9 with =7.5 measured by Koch et al. (2006). This confirms that our stars are galaxy members. The spectra were normalized using DAOSPEC (Stetson & Pancino, 2008) for each of the wavelength ranges presented in Table 1. We used a 20 to 40 degree polynomial fit.
@@ 3.1 Line list and model atmospheres
We adopted the new MARCS 1D atmosphere models and selected the Standard composition class, that is, we included the classical -enhancement of +0.4 dex at low metallicity. They were downloaded from the MARCS web site (Gustafsson et al., 2008), and interpolated using Thomas Masseron’s code, which is available on the same web site.222http://marcs.astro.uu.se Inside a cube of eight reference models, this code performs a linear interpolation on three given parameters: Teff, g, and [Fe/H].
@@ 3.2 Photometric temperature and gravity
The atmospheric parameters were initially determined using photometric information, as reported in Table 2. The first approximated determination of the stellar effective temperature (Teff) was based on the VI, VJ, VH, and VK color indices measured by Battaglia et al. (2006), and the J and Ks photometry was taken from the VISTA commissioning data, which were also calibrated onto the 2MASS photometric system. We assumed (Cardelli et al., 1989) and = 0.03 for Fornax (Letarte et al., 2010), and = 0.061 for Carina (de Boer et al., 2014) for the reddening correction. The adopted photometric effective temperatures () are listed in Table 3. They correspond to the simple average of the four color temperatures derived from , , , and with the calibration of Ramírez & Meléndez (2005).
@@
Because very few Fe ii lines can be detected in our X-Shooter spectra, the determination of surface gravities from the ionization balance of Fe I versus Fe II was not possible. Non-local thermodynamic equilibrium (NLTE) effects also play a role at extremely low metallicity, and impact the abundances of Fe I, with (Fe II–Fe I) up to +0.20 dex at [Fe/H] = (Mashonkina et al., 2017a), and thus surface gravities were determined from their relation with ,
@@
assuming = 4.44, = 5790 K, and Mbol⊙ = 4.75 for the Sun. We adopted a stellar mass of 0.8 M⊙ and calculated the bolometric corrections using the Alonso et al. (1999) calibration, with a distance of d=138 kpc (Battaglia et al., 2006) for Fornax and d=106 kpc (de Boer et al., 2014) for Carina.
3.3 Final stellar parameters and abundance determinations